# Breaking the Ceiling of the LLM Community by Treating Token Generation as a Classification for Ensembling

### v1

@@ 1 Introduction
Taking computer vision (CV) classification as an example, it is common to ensemble the output probability vectors of multiple models (e.g. by averaging) to achieve superior results (Krizhevsky et al., 2017). This approach remains effective even with recent CV models. As shown in Tab.1, we selected several common CV models (Chen et al., 2022; Tan and Le, 2019; Wang et al., 2022) for ensembling and observed better accuracy on ImageNet (Deng et al., 2009) compared to using a single model. Similarly, the popular decoder-only LLM architecture generates text by producing tokens one by one, with each generation step resulting in a probability vector of the length of the vocabulary. Inspired by CV, we propose to treat each generation step as a classification task, and by ensembling multiple models, we can achieve higher accuracy, as shown in Fig.1. There is already work that simplifies problems into binary tasks, exploiting the collective wisdom of LLMs and achieving better results (Schoenegger et al., 2024), demonstrating the feasibility of this approach.
@@
Another advantage is that early errors in LLMs often snowball into later errors (Zhang et al., 2023). Ensembling during generation helps prevent the generation of inaccurate tokens at each step, thereby reducing misleading cues for subsequent token generation. In this paper, we conducted experiments at several points in time between November 2023 and June 2024, ensembling available state-of-the-art (SOTA) LLMs up to each of these points. We found that this approach significantly outperformed any single model available at those times on five popular benchmarks involving subject examination, mathematics, reasoning, and knowledge-based QA.
In addition, we found that for text generation it seemed unnecessary to ensemble every token. For example, for the question "What Andean animal has banana-shaped ears?" shown in Fig.1, the most critical part is for the LLM to generate the key token "llama". The initial part of the answer "It should be _" or "The animal is _" do not significantly affect the correctness of the final answer. Ideally, the step that produces the token "llama" is the one we want to ensemble.
@@ 3.1 Overall Process of GaC
When generating text, LLMs output a probability vector of the same dimension as their vocabulary. Given $n$ LLMs to be ensembled, we first take the union of their vocabularies and create a mapping matrix that can project the probability vectors to the union dimensions (Sec.3.2). At each generation step, all LLMs produce outputs that are mapped to the union vocabulary dimensions and ensembled to sample the next token. The tokenizer of each LLM then converts the sampled token into token IDs for the next step (Sec.3.3). As mentioned in Sec.1, not all tokens have the necessity for ensembling, so we also try to ensemble only certain key tokens (Sec.3.4).
@@ 3.2 Creating the Union Mapping
During this process, we record the positions of tokens from $V^{i}$ in $V^{U}$ and create corresponding mapping matrices $\mathbf{M}^{i}\in\mathbb{R}^{|V^{i}|\times|V^{U}|}$.
@@ 3.3 GaC Ensembling
We repeat (3) and (4) until the stopping criteria are met, such as outputting an end-of-sentence token or reaching the maximum length, as shown in Fig.3. In our implementation, different LLMs run in parallel on different GPUs, so the duration of each step is equal to the time taken by the slowest LLM. Since we have not modified a complete forward pass, our approach is compatible with techniques such as vLLM, DeepSpeed, quantization, and hardware optimizations (Kwon et al., 2023; Rasley et al., 2020).
@@ 3.4 Ensembling Key Tokens with Threshold
As mentioned in the last part of Sec.1, most tokens do not significantly affect the correctness of the response. From Tab.1, we can see that LLMs and CV models have similar ECE levels, suggesting that the confidence scores of LLMs may reflect accuracy to some extent. Therefore, we also experiment with ensembling only the tokens with confidence below a threshold $t$. We choose a model as the gate, denoted $\textit{LLM}_{g}$, and use its maximum probability at each step as the confidence score. During the ensemble, we replace the original (3) with:
@@ 4.1 Overview
In this section, we first present the experimental setup, including the benchmarks and hardware used (Sec.4.2). We then test the effects of different ensemble weights for GaC (Sec.4.3) and compare it with other methods (Sec.4.4). We also select SOTA LLMs available at different times for ensembling to explore the performance ceiling at each time period (Sec.4.5). Finally, we experiment with thresholded ensembling to explore variations in latency and performance (Sec.4.6).
@@ 4.2 Experimental Settings
Benchmarks. GaC is not limited to specific tasks, so we tested it as broadly as possible. We selected a total of five benchmarks. For general capabilities, we chose MMLU (Hendrycks et al., 2020). For maths, we utilized GSM8K (Cobbe et al., 2021). For reasoning, we employed BBH (Suzgun et al., 2023). For knowledge capabilities, we included TriviaQA (Joshi et al., 2017) and NaturalQuestions (NQ) (Kwiatkowski et al., 2019). Note that all scores, including those for individual models, were computed locally under the same environment to ensure fairness [^2] Please see Appendix.A for more benchmarks details., using lm-evaluation-harness [^3] https://github.com/EleutherAI/lm-evaluation-harness v0.4.1 Gao et al. (2023).
Hardware and Latency. Each LLM was loaded on 1(n) A100 GPU(s) according to its memory requirements [^4] We listed each model and its hardware in Appendix.B., using naive model parallelism (HuggingFace, 2024) without optimization for inference. During ensembling, different LLMs were loaded on separate GPU(s) and executed in parallel, managed and communicated via Ray (Moritz et al., 2018). We also recorded the latency (ms/token). Each model performs a "dry run" after being loaded onto the GPU, generating 1024 tokens to warm up CUDA before experimentation, following the practice of Mehta et al. (2024).
4.3 Different Ensemble Weights
@@
Before proceeding with further experiments, we tested different ensemble weights for GaC. We used a simple averaging of the probabilities from each model in Eq.3. We now replace Eq.3 by $\frac{1}{\sum_{i}w^{i}}\sum_{i}w^{i}p^{i}(\cdot\,|\,\mathcal{I}^{\,i})\mathbf{M}^{i}$, where $w^{i}$ is the ensemble weight for $\textit{LLM}_{i}$. We set $w^{i}$ separately to each LLM’s score on MMLU, the MMLU score minus the ECE and 1 (i.e. averaging). We selected two sets of LLMs (Wang et al., 2023a; Kim et al., 2023; Young et al., 2024; Bai et al., 2023; NousResearch, 2024a) listed in Tab.2 with different sizes for GaC ensemble. The results on MMLU are shown in Fig.4. We observed no significant differences between the different weights, so for simplicity we decided to use averaging.
@@ 4.4 Comparison with Other Methods
Table 3: Results of comparison with other methods. Upper and lower halves represent different ensemble combinations. Blue indicates the best result for each ensemble. For id 8, GaC bottom right shows ensembled token proportion, top right shows threshold.
@@ 4.5 Breaking the Ceiling
However, with the release of Qwen2-72B-Instruct, which showed comparable performance to Llama-3-70B-Instruct, the ensemble again led to significant improvements (row id 10). In rows 11 and 12, we ensemble the top two best-performing models for each benchmark at the two most recent times, including the challenging time of 2024/04/18, and observe performance gains with this task-specific top-two ensemble even on 04/18 (row id 11). Finally, row id 12 shows the best results available for the open source community on 2024/06/07. By pushing the boundaries of the community, we can narrow the gap with proprietary models and promote the democratization of LLMs.
4.6 Ensemble with Threshold
@@ 5 Conclusion
Appendix A Benchmarks
@@
Appendix B Hardware Specifications

### v2

@@ 1 Introduction
Taking computer vision (CV) classification as an example, it is common to ensemble the output probability vectors of multiple models (e.g. by averaging) to achieve superior results (Krizhevsky et al., 2017). This approach remains effective even with recent CV models. As shown in Tab.1, we selected several common CV models (Chen et al., 2022; Tan and Le, 2019; Wang et al., 2022) for ensembling and observed better accuracy on ImageNet (Deng et al., 2009) compared to using a single model. Similarly, the popular decoder-only LLM architecture generates text by producing tokens one by one, with each generation step resulting in a probability vector of the length of the vocabulary. Inspired by CV, we propose to treat each Generation step as a Classification task (GaC), and by ensembling multiple models, we can achieve higher accuracy, as shown in Fig.1. There is already work that simplifies problems into binary tasks, exploiting the collective wisdom of LLMs and achieving better results (Schoenegger et al., 2024), demonstrating the feasibility of this approach.
@@
Another advantage is that early errors in LLMs often snowball into later errors (Zhang et al., 2023). Ensembling during generation helps prevent the generation of inaccurate tokens at each step, thereby reducing misleading cues for subsequent token generation. In this paper, we conducted experiments at several points in time between November 2023 and June 2024, ensembling available state-of-the-art (SOTA) LLMs up to each of these points. We found that this approach significantly outperformed any single model available at those times as well as existing methods on five popular benchmarks involving subject examination, mathematics, reasoning, and knowledge-based QA.
In addition, we found that for text generation it seemed unnecessary to ensemble at every step. For example, for the question "What Andean animal has banana-shaped ears?" shown in Fig.1, the most critical part is for the LLM to generate the key token "llama". The initial part of the answer "It should be _" or "The animal is _" do not significantly affect the correctness of the final answer. Ideally, the step that produces the token "llama" is the one we want to ensemble.
@@
In summary, we propose a framework GaC, which allows multiple heterogeneous LLMs to jointly decide the next generated token during text generation:
  • When we ensemble at every generation step, experiments show that it outperforms any single SOTA LLM from different time periods and existing ensemble methods.
  • By using cascade to ensemble only at important steps during text generation, experiments show better results with lower latency.
@@ 3.1 Overall Process of GaC
When generating text, LLMs output a probability vector of the same dimension as their vocabulary. Given $n$ LLMs to be ensembled, we first take the union of their vocabularies and create a mapping matrix that can project the probability vectors to the union dimensions (Sec.3.2). At each generation step, all LLMs produce outputs that are mapped to the union vocabulary dimensions and ensembled to sample the next token. The tokenizer of each LLM then converts the sampled token into token IDs for the next step (Sec.3.3). As mentioned in Sec.1, not all generation steps have the necessity for ensembling, so we also try to ensemble only certain key tokens (Sec.3.4).
@@ 3.2 Creating the Union Mapping
During this process, we record the positions of tokens from $V^{i}$ in $V^{U}$ and create corresponding mapping matrices $\mathbf{M}^{i}\in\{0,1\}^{|V^{i}|\times|V^{U}|}$.
@@ 3.3 GaC Ensembling
We repeat (3) and (4) until the stopping criteria are met, such as outputting an end-of-sentence token or reaching the maximum length, as shown in Fig.3.222We provide step-by-step examples in Appx.A. In our implementation, different LLMs run in parallel on different GPUs, so the duration of each step is equal to the time taken by the slowest LLM. Since we have not modified a complete forward pass, our approach is compatible with techniques such as vLLM, DeepSpeed, quantization, and hardware optimizations (Kwon et al., 2023; Rasley et al., 2020).
@@ 3.4 Ensembling Key Tokens with Threshold
As mentioned in the last part of Sec.1, most tokens do not significantly affect the correctness of the response. From Tab.1, we can see that LLMs and CV models have similar ECE levels, suggesting that the confidence scores of LLMs may reflect accuracy to some extent. Therefore, we also experiment with ensembling only the steps with confidence below a threshold $t$. We choose a model as the gate, denoted $\textit{LLM}_{g}$, and use its maximum probability at each step as the confidence score. During the ensemble, we replace the original (3) with:
@@ 4.1 Overview
In this section, we first present the experimental setup, including the benchmarks and hardware used (Sec.4.2). We then test the effects of different ensemble parameters for GaC (Sec.4.3) and compare it with other methods (Sec.4.4). We also select SOTA LLMs available at different times for ensembling to explore the performance ceiling at each time period (Sec.4.5). Finally, we experiment with thresholded ensembling to explore variations in latency and performance (Sec.4.6).
@@ 4.2 Experimental Settings
Benchmarks. GaC is not limited to specific tasks, so we tested it as broadly as possible. We selected a total of five benchmarks. For general capabilities, we chose MMLU (Hendrycks et al., 2020). For maths, we utilized GSM8K (Cobbe et al., 2021). For reasoning, we employed BBH (Suzgun et al., 2023). For knowledge capabilities, we included TriviaQA (Joshi et al., 2017) and NaturalQuestions (NQ) (Kwiatkowski et al., 2019). Note that all scores, including those for individual models, were computed locally under the same environment to ensure fairness [^3] Please see Appx.B for more benchmarks details., using lm-evaluation-harness [^4] https://github.com/EleutherAI/lm-evaluation-harness v0.4.1 Gao et al. (2023).
Hardware and Latency. Each LLM was loaded on 1(n) A100 GPU(s) according to its memory requirements [^5] We listed each model and its hardware in Appx.C., using naive model parallelism (HuggingFace, 2024) without optimization for inference. During ensembling, different LLMs were loaded on separate GPU(s) and executed in parallel, managed and communicated via Ray (Moritz et al., 2018). We also recorded the latency (ms/token). Each model performs a "dry run" after being loaded onto the GPU, generating 1024 tokens to warm up CUDA before experimentation, following the practice of Mehta et al. (2024).
4.3 Exploring Ensemble Parameters
@@
Different Ensemble Weights. Before proceeding with further experiments, we tested different ensemble weights for GaC. We used a simple averaging of the probabilities from each model in Eq.3. We now replace Eq.3 by $\frac{1}{\sum_{i}w^{i}}\sum_{i}w^{i}p^{i}(\cdot\,|\,\mathcal{I}^{\,i})\mathbf{M}^{i}$, where $w^{i}$ is the ensemble weight for $\textit{LLM}_{i}$. We set $w^{i}$ separately to each LLM’s score on MMLU, the MMLU score minus the ECE and 1 (i.e. averaging). We selected two sets of LLMs (Wang et al., 2023a; Kim et al., 2023; Young et al., 2024; Bai et al., 2023; NousResearch, 2024a) listed in Tab.2 with different model sizes for GaC ensemble. The results on MMLU are shown in Fig.4. We observed no significant differences between the different weights, so for simplicity we decided to use averaging.
Model Acc [%] $\Delta$
Qwen1.5-14B-Chat 67.20 -
Phi-3-mini-4k-instruct 67.11 -
OpenChat-3.5-0106 64.53 -
SOLAR-10.7B-Instruct-v1.0 64.48 -
GaC (Qwen1.5 + Phi-3) 69.91 +2.71
GaC (Qwen1.5 + Phi-3 + OpenChat) 70.40 +3.20
GaC (Qwen1.5 + Phi-3 + OpenChat
+ SOLAR) 70.01 +2.81
Table 3: Performance of different ensemble sizes on MMLU. The top part lists individual models, while the bottom part shows ensemble results (model names abbreviated). $\Delta$ represents the improvement compared to the best individual model.
Different Ensemble Sizes. We also investigated the effect of ensembling different numbers of models on performance. Specifically, we selected four popular small LLMs (Bai et al., 2023; Kim et al., 2023; Wang et al., 2023a; Abdin et al., 2024). After sorting the models based on their MMLU scores from highest to lowest, we sequentially ensembled the top 2, 3, and 4 models. As shown in Tab.3, the results indicate that the performance gain from ensembling does not always increase with the addition of more models. The highest score was achieved when ensembling three models, while ensembling all four models resulted in a slight decrease in performance (from 70.40 to 70.01).
@@ 4.4 Comparison with Other Methods
Table 4: Results of comparison with other methods. Upper and lower halves represent different ensemble combinations. Blue indicates the best result for each ensemble. For id 8, GaC bottom right shows ensembled token proportion, top right shows threshold. For id 6 and 14, we ensemble at every generation step.
@@ 4.5 Breaking the Ceiling
However, with the release of Qwen2-72B-Instruct, which showed comparable performance to Llama-3-70B-Instruct, the ensemble again led to significant improvements (row id 10). In rows 11 and 12, we ensemble the top two best-performing models for each benchmark at the two most recent times, including the challenging time of 2024/04/18, and observe performance gains with this task-specific top-two ensemble even on 04/18 (row id 11). Finally, row id 12 shows the best results available for the open source community on 2024/06/07. By pushing the boundaries of the community, we can narrow the gap with proprietary models and promote the democratization of LLMs. [^8] We also tested translation and summarization in Appx.D.
@@ 5 Conclusion
Appendix A Step-by-Step Examples of GaC
To better demonstrate the GaC ensemble, we posed two classic and interesting questions to two individual models (OpenChat-3.5-0106 and SOLAR-10.7B-Instruct-v1.0), as well as to their ensemble.
[…]
  • Q2: 9.11 and 9.9, which is bigger?
Their responses are listed in Tab.8. Next, we list the tokens generated by the two models at each step during the ensemble process that have the highest confidence, as well as the token chosen by the GaC, in Tab.9.
In Q1, OpenChat tried to output incorrect answers "ste" and "ty" directly in step 1 and step 4, but they were not adopted due to insufficient confidence. In step 4, SOLAR outputted the high confidence token "car", which resulted in the ensemble answer being correct.
Response of Q1 Response of Q2
OpenChat-3.5-0106 steering wheel. 9.11 is bigger than 9.9.
SOLAR-10.7B-Instruct-v1.0 The word "car" does not belong with the others as the other options are components or parts of a car, while "car" is the overall vehicle itself. To determine which number is bigger, we compare the numbers digit by digit from the rightmost digit (the decimal point) to the left. In this case, we have 1 in the decimal place for 9.11 and 9 in the decimal place for 9.9. Since 9 is greater than 1, we can conclude that 9.9 is the bigger number.
GaC (OpenChat + SOLAR) The word "car" does not belong with the others. The other words (tyre, steering wheel, and engine) are all parts of a car, while "car" is the whole vehicle itself. 9.9 is bigger than 9.11. However, it seems like there might be a mistake in the numbers provided. The number 9.11 is commonly associated with the date of the September 11, 2001 terrorist attacks in the United States, and it is not a typical number to compare with 9.9. If you meant to compare 9.1 and 9.9, then 9.9 is indeed bigger.
Ground truth Car does not belong because all others are components of a car. 9.9
Table 8: Responses and groud truth of Q1 and Q2
Step Q1 Ensemble (OpenChat / SOLAR / GaC) Q2 Ensemble (OpenChat / SOLAR / GaC)
1 ste (0.419) / The (0.950) / The   (0.976) / To (0.603) /
[…]
3  that (0.594) /  " (0.999) /  " 9 (0.942) / When (0.342) / 9
4 ty (0.714) / car (1.000) / car . (1.000) / . (1.000) / .
5 " (0.999) / " (1.000) / " 1 (0.689) / 9 (1.000) / 9
6  does (0.992) /  does (0.998) /  does  is (0.994) /  is (1.000) /  is
7  not (1.000) /  not (1.000) /  not  bigger (0.969) /  a (0.765) /  bigger
8  belong (0.999) /  belong (0.950) /  belong  than (0.645) /  than (1.000) /  than
9  with (0.999) /  with (0.990) /  with   (1.000) /   (1.000) /
10  the (1.000) /  the (1.000) /  the 9 (1.000) / 9 (1.000) / 9
11  others (0.889) /  others (0.990) /  others . (1.000) / . (1.000) / .
12 . (0.726) / as (0.638) / . 1 (1.000) / 1 (1.000) / 1
13  The (0.428) /  The (0.942) /  The 1 (1.000) / 1 (1.000) / 1
14  other (0.866) /  other (0.982) /  other . (0.948) / . (0.483) / .
Table 9: The highest confidence token in each step of individual models (confidence indicated in parentheses) and the token chosen by GaC.
Appendix B Benchmarks
@@
Appendix C Hardware Specifications
We list the hardware corresponding to the models used in this paper in Tab.10.
@@ Appendix C Hardware Specifications
Appendix D Translation and Summarization
We applied the GaC ensemble to OpenChat-3.5-0106 and SOLAR-10.7B-Instruct-v1.0 to evaluate its performance on translation and summarization tasks (Bojar et al., 2014, 2016; Narayan et al., 2018). The results are presented in Tab.11.
Models EN → FR FR → EN EN → DE DE → EN XSUM
BLEU CHRF BLEU CHRF BLEU CHRF BLEU CHRF ROUGE-L
OpenChat-3.5-0106 35.40 61.71 40.52 64.82 31.48 59.84 44.32 67.97 26.54
SOLAR-10.7B-Instruct-v1.0 35.61 62.15 40.28 64.86 30.90 60.06 42.85 67.79 23.81
GaC (OpenChat + SOLAR) 36.59 62.82 41.63 65.56 33.10 61.43 44.65 68.69 27.26
$\Delta$ +0.98 +0.67 +1.11 +0.70 +1.62 +1.37 +0.33 +0.72 +0.72
Table 11: Performance on translation and summarization tasks. $\Delta$ represents the improvement compared to the best individual model.